OpenStudy (anonymous):

I have a problem that just gives me a graph and it asks for the following things. (f+g) (f-g) (fg) (f/g) There is not any equations or anything. Just points on the graph that make up the two lines. All of the questions like this that I have done always have something like f(x)=x^2-1. This question just asks what f-g, and etc. No equations just points. How do I find the answer with just the points?

OpenStudy (anonymous):

Here is a photo.

OpenStudy (anonymous):

Glancing at it it would appear you find the value they are asking and apply the function. For example: Problem 49. @x = -3 \[f(x) = 4\] \[g(x)= 1\] \[(f+g) = 4+1=5\]

OpenStudy (anonymous):

The graphing question is just applying the math to each of the points and drawing the new graph.

OpenStudy (anonymous):

Yes I don't know how to do that. There isn't any equations in which I can do that.

OpenStudy (anonymous):

You have it evaluate it one point at at time. We have evaluated x=-3 for example for (f+g) and got 5. What is x=-2?

OpenStudy (anonymous):

What it looks like to me is. Here are some values of x and y congratulations you should be able to find something from this. I'm sorry we did make for time to actually explain this concept to you. Good luck on your teacher's test.

OpenStudy (anonymous):

Those are easy ones. You just find out what the out put of x=-3 is on the graph and put it in the equation. I know how to do that. I'm talking about the one that is Find the domain of f+g

OpenStudy (anonymous):

f(x) stops having values along the x axis at f(3). That is why the upper range of (f+g) stops there. There is no more values of f(x) to add in. The lower end both go to x=-4. Again f(x) stops there even though g(x) has a value for -5.

OpenStudy (anonymous):

*There are no more values..... Bad typos....

OpenStudy (anonymous):

I don't care about that problem. I have already stated that I can do those when the book say what is f(x)+g(x) when x=-3 these are easy. Just look for the output. The other question asked about find the domain of f+g or all of f plus all of g..

OpenStudy (anonymous):

On the divsion. The reason you have a less than sign and not an equals is g(-4) and g(3) are both 0 which when in the denominator is undefined.

OpenStudy (anonymous):

I just explained the domain question. Relook at the 3 posts prior, to include a typo fix. The same principles apply.

OpenStudy (anonymous):

I don't care about that problem. I have already stated that I can do those when the book say what is f(x)+g(x) when x=-3 these are easy. Just look for the output. The other question asked about find the domain of f+g or all of f plus all of g..

OpenStudy (anonymous):

f+g x=?

OpenStudy (anonymous):

I don't care about that problem. I have already stated that I can do those when the book say what is f(x)+g(x) when x=-3 these are easy. Just look for the output. The other question asked about find the domain of f+g or all of f plus all of g..

OpenStudy (anonymous):

55

OpenStudy (anonymous):

Your picture has the answer right beside it in blue.... I was explaining why that answer was right. I assume you wanted to know why the book answer was correct.

OpenStudy (anonymous):

I don't care about that problem. I have already stated that I can do those when the book say what is f(x)+g(x) when x=-3 these are easy. Just look for the output. The other question asked about find the domain of f+g or all of f plus all of g..

OpenStudy (anonymous):

No. I know the book is correct. I want to know how to do seemingly impossible problems like 55 and 56

OpenStudy (anonymous):

I don't care about that problem. I have already stated that I can do those when the book say what is f(x)+g(x) when x=-3 these are easy. Just look for the output. The other question asked about find the domain of f+g or all of f plus all of g..

OpenStudy (anonymous):

And why in problem 53 is -5 and 5 not included

OpenStudy (anonymous):

Since I have explained why the domain is limited and it didn't seem clear perhaps you should try chat and see if someone else can explain it to you. It's pssoibel I'm not epxlaining it in a way that you can understand.

OpenStudy (anonymous):

I don't care about that problem. I have already stated that I can do those when the book say what is f(x)+g(x) when x=-3 these are easy. Just look for the output. The other question asked about find the domain of f+g or all of f plus all of g..

OpenStudy (anonymous):

Looks like random guess work. The concepts in my book are not defined well. They just say here you go here is the answer. By the way here are some more problems that are harder that we are not going to explain. Have fun

OpenStudy (anonymous):

I don't care about that problem. I have already stated that I can do those when the book say what is f(x)+g(x) when x=-3 these are easy. Just look for the output. The other question asked about find the domain of f+g or all of f plus all of g..

OpenStudy (anonymous):

Looks like random guess work. The concepts in my book are not defined well. They just say here you go here is the answer. By the way here are some more problems that are harder that we are not going to explain. Have fun

OpenStudy (anonymous):

Osirisis were you basically saying for adding take all the terms X in f and g and add the ranges?